Calculate the pH at 25°C of a 0.76M solution of anilinium chloride C6H5NH3Cl. Note that aniline C6H5NH2 is a weak base with a pKb of 4.87.
Answer #1

Answer

pH = 4.63

Explanation

Conjucate acid is partly dissociates in water

C6H5NH3+(aq) + H2O(l) -------> C6H5NH2(aq) + H3O+(aq)

Ka = [C6H5NH2][H3O+]/[C6H5NH3+]

Ka = Kw/Kb = 1.00 ×10-14/1.35×10-5= 7.41× 10-10

at equillibrium

[C6H5NH3+] = 0.76 - x

[C6H5NH2] = x

[H3O+] = x

so,

x2/(0.76 - x) = 7.41×10-10

we can assume , 0.76 -x = 0.76 because x is small value

x2/0.76 = 7.41 ×10-10

x2 = 5.63× 10-10

x = 2.37×10-5

[H3O+] = 2.37×10-5M

pH = -log[H3O+]

pH = -log(2.37 ×10-5)

pH = 4.63